An apartment has the dimensions 17 m by 9 m by 6 m. The temperature is 20°C, and the relative humidity is 50 percent. What is the total mass (in kg) of water vapor in the air in the apartment?
1
Expert's answer
2019-04-22T11:14:05-0400

For this temperature - 20 C the maximal humidity is

"\rho_m=17.3 (\frac{g}{m^3})""\rho_a=0.5*17.3=8.65 (\frac{g}{m^3})"

for mass water vapor

"m=17\cdot 9\cdot 6\cdot 8.65 =7.94 kg"